40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 7 17:44:10 2014, diego, Update, LSC, IR Resonance Script Status wavelets.nb.zip
    Reply  Wed Dec 3 01:18:15 2014, diego, Update, LSC, IR Resonance Script Status 
       Reply  Wed Dec 3 02:01:57 2014, Koji, Update, LSC, IR Resonance Script Status 
Message ID: 10749     Entry time: Wed Dec 3 02:01:57 2014     In reply to: 10747
Author: Koji 
Type: Update 
Category: LSC 
Subject: IR Resonance Script Status 

The other night (before the holidays), I tried ALS offset tuning  with IR POX/POY signals and it worked pretty good.
I didn't need to tune the offset after the scanning script stopped.

Once we are at the foot hill of the main resonance, I ran something like

ezcaservo -r C1:LSC-POX11_I_MON C1:LSC-ALSX_OFFSET -g -0.003 &
ezcaservo -r C1:LSC-POY11_I_MON C1:LSC-ALSY_OFFSET -g -0.003 &

(... I am writing this with my memory. I could be wrong but conceptually the commands looked like these)

ELOG V3.1.3-